# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
where $F_{1}$ is as in the proof. The second isomorphism follows from the Quillen adjunction in 5.5(ii) and the third isomorphism holds by 5.5(i) because $\chi\neq\chi^{\prime}$ for any other face in $\mathscr{F}_{0}$. This tells us that $\dim_{k}[\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}}=1$ in this case.
This could suggest that the functor ${\operatorname{Ho}}(\Delta)$ might be faithful in general, perhaps identifying bijectively a morphism set $[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}_{\operatorname{aff}}}}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ with the set of all $(f_{F})\in\prod_{F\subseteq\overline{C}}[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}$ such that ${\operatorname{Ho}}({\operatorname{Res}}_{{\mathcal{H}}_{F}}^{{\mathcal{H}}_{F^{\prime}}})(f_{F^{\prime}})=f_{F}$ whenever $F^{\prime}\subseteq\overline{F}$. When the group $\mathbb{G}$ has semisimple rank 1, this is in fact true. Indeed, as the complex (5.3) is a short exact sequence in that case and since ${\mathcal{H}}_{C}$ is always a semisimple algebra, we have that the augmentation again realises $\bigoplus_{F\in\mathscr{F}_{0}}{\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F}}{\mathfrak{m}}$ as a Gorenstein projective replacement of ${\mathfrak{m}}$, for any ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$. After chasing through the adjunctions one deduces, for any ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$, that the isomorphism $[{\mathfrak{m}},\mathfrak{n}]_{\mathcal{H}_{\operatorname{aff}}}\cong\bigoplus_{F\in\mathscr{F}_{0}}[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}$ obtained from the augmentation agrees with the map induced by ${\operatorname{Ho}}(\Delta)$. Finally, the compatibility condition between the $f_{F}$’s with $F\in\mathscr{F}_{0}$ and $f_{C}$ is empty here because ${\operatorname{Ho}}({\mathcal{H}}_{C})=0$.
As an application of 5.7, we deduce:
@@
Suppose that ${\mathfrak{m}},{\mathfrak{m}}^{\prime}\in{\operatorname{Mod}}({\mathcal{H}})$ are two simple supersingular modules, parametrised by the pairs $(\chi,V)$ and $(\chi^{\prime},V^{\prime})$ respectively, such that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$. Assume that ${\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}})=\infty$ and that $\chi,\chi^{\prime}$ and $\Phi$ are not as in 5.7(i). Then $\chi^{\prime}=\chi^{\tilde{\omega}}$ for some $\tilde{\omega}\in\widetilde{\Omega}$.
@@
By 2.13, we have that restriction of scalars ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}_{\operatorname{aff}}}}$ is a right Quillen functor which preserves all weak equivalences. If ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ we therefore have that ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}_{\operatorname{aff}}}})({\mathfrak{m}})$ and ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}_{\operatorname{aff}}}})({\mathfrak{m}}^{\prime})$ are also isomorphic in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. By the proof of [32, Proposition 6.17] we have
@@
Assume that $\chi^{\prime}\neq\chi^{\tilde{\omega}}$ for any $\tilde{\omega}\in\widetilde{\Omega}$. Since we get from the above that
@@
it then follows from 5.7(ii) that $[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{\mathcal{H}_{\operatorname{aff}}}$ has image zero under ${\operatorname{Ho}}(\Delta)$. If $f:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ denotes an isomorphism in ${\operatorname{Ho}}({\mathcal{H}})$, then this implies that ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}})(f)$ is both zero and an isomorphism for every $F\subseteq\overline{C}$. We deduce that ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$, and thus ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$, is projective for every $F\subseteq\overline{C}$. Applying 5.5(iii) and 5.4 it follows that ${\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}})<\infty$, contradicting our assumptions.
The conclusion of 5.17 is false when we are in the setting of 5.7(i), as is visible from the following examples.
@@
Suppose that $G=\mathrm{PGL}_{3}(\mathfrak{F})$ and assume that $p\neq 3$. Then $\Omega=\langle\omega\rangle\cong{\mathbb{Z}}/3{\mathbb{Z}}$ where $\omega$ is the image of $\left(\begin{smallmatrix}0&0&\varpi\\ 1&0&0\\ 0&1&0\end{smallmatrix}\right)$ in $W$. Similarly, let $\hat{\omega}$ denote the image of the same matrix in $\widetilde{W}$. We have $S=\{s,s^{\prime},s^{\prime\prime}\}$ and let $\chi\neq\chi^{\prime}$ be two supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ as in 5.7(i), i.e. corresponding to the pairs $(J,\xi)$ and $(J^{\prime},\xi)$ with $S_{\xi}=S$, $J=\{s,s^{\prime\prime}\}$ and $J^{\prime}=\{s^{\prime\prime}\}$. We will use the notation $\chi=(-1,0,-1)_{\xi}$ and $\chi^{\prime}=(0,0,-1)_{\xi}$ for these characters, the entries of the tuple corresponding to the value of the character at $T_{\hat{s}}$, $T_{\hat{s^{\prime}}}$ and $T_{\hat{s^{\prime\prime}}}$, taken in that order. Note that $S=S_{\xi}$ forces the character $\xi$ to either be trivial or to be $\det^{m}$ when there exists $m\in{\mathbb{N}}$ such that $q=3m+1$, and in particular it is fixed by $\omega$ (c.f. 5.26 later). Also, the element $\omega$ acts on the set $S$ by permuting its elements cyclically and correspondingly we see that $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}=T(\mathbb{F}_{q})$ and ${\mathcal{H}}_{\chi}={\mathcal{H}}_{\chi^{\prime}}={\mathcal{H}_{\operatorname{aff}}}$. If we let $V$ and $V^{\prime}$ be two irreducible finite dimensional representations of $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$ and $\chi^{\prime}$ respectively, we then automatically have $V=\chi$ and $V^{\prime}=\chi^{\prime}$. Thus we obtain two nonisomorphic simple supersingular ${\mathcal{H}}$-modules by taking ${\mathfrak{m}}=\chi\otimes_{{\mathcal{H}_{\operatorname{aff}}}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=\chi^{\prime}\otimes_{{\mathcal{H}_{\operatorname{aff}}}}{\mathcal{H}}$. Furthermore, those have infinite projective dimension (c.f. 5.4). We claim that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$.
To see this, first note that
(5.6) ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}{\mathfrak{m}}\cong\chi_{0}\oplus\chi_{1}\oplus\chi_{2}=(-1,0,-1)_{\xi}\oplus(-1,-1,0)_{\xi}\oplus(0,-1,-1)_{\xi}$
and similarly
(5.7) ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}{\mathfrak{m}}^{\prime}\cong\chi^{\prime}_{0}\oplus\chi^{\prime}_{1}\oplus\chi^{\prime}_{2}=(0,0,-1)_{\xi}\oplus(-1,0,0)_{\xi}\oplus(0,-1,0)_{\xi}.$
@@
Next, observe that for every face $F\subseteq\overline{C}$ with $F\neq C$ the stabiliser $\Omega_{F}$ of $F$ in $\Omega$ is trivial. Thus, for such faces, ${\mathcal{H}}^{\dagger}_{F}={\mathcal{H}}_{F}$ by [25, Lemma 4.20(i)]. Meanwhile, ${\mathcal{H}}_{C}^{\dagger}\cong k[\widetilde{\Omega}]$ is the group algebra of a finite group. As $\left|\widetilde{\Omega}\right|=3\left|T(\mathbb{F}_{q})\right|$ is not divisible by $p$, this algebra is semisimple and thus ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{C}^{\dagger}}{\mathfrak{m}}$ is projective. Let $x$ and $e$ denote respectively the vertex and edge of $\overline{C}$ with $S_{x}=\{s^{\prime},s^{\prime\prime}\}$ and $S_{e}=\{s^{\prime}\}$. Since all faces in $\overline{C}$ of a given dimension are $\Omega$-conjugate to each other, the Ollivier-Schneider resolution of ${\mathfrak{m}}$ (c.f.  (5.1)) is given by
(5.8) $0\to{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{C}^{\dagger}}({\mathfrak{m}})\otimes_{{\mathcal{H}}_{C}^{\dagger}}(\epsilon_{C}){\mathcal{H}}\to{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{e}}({\mathfrak{m}})\otimes_{{\mathcal{H}}_{e}}(\epsilon_{e}){\mathcal{H}}\to{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\otimes_{{\mathcal{H}}_{x}}{\mathcal{H}}\stackrel{{\scriptstyle\varepsilon}}{{\to}}{\mathfrak{m}}\to 0,$
where we drop the twist by orientation $\epsilon_{x}$ character from the notation as it is trivial. By 5.10(ii) we see that, for each $0\leq i\leq 2$, the character ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{e}}(\chi_{i})$ is projective. Thus (5.8) tells us that $\varepsilon:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\otimes_{{\mathcal{H}}_{x}}{\mathcal{H}}\to{\mathfrak{m}}$ is a trivial fibration. Completely analogously, $\varepsilon:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}}^{\prime})\otimes_{{\mathcal{H}}_{x}}{\mathcal{H}}\to{\mathfrak{m}}^{\prime}$ is also a trivial fibration.
@@
Now, by appealing to 5.10(ii) we see that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{2})$ and ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{1})$ are projective, while ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{0})={\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{0})$ and ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{1})={\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{2})$ are not projective. By plugging all this into (5.6) and (5.7), we get a sequence of isomorphisms
${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\cong{\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{0})\oplus{\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{1})={\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{0})\oplus{\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{2})\cong{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}}^{\prime})$
@@
Suppose that $G=\mathrm{GL}_{3}(\mathfrak{F})$. Then $\Omega=\langle\omega\rangle\cong{\mathbb{Z}}$ where $\omega$ is once again the image of $\left(\begin{smallmatrix}0&0&\varpi\\ 1&0&0\\ 0&1&0\end{smallmatrix}\right)$ in $W$. We keep the same notation as in the previous example, so we let $\hat{\omega}$ be the same lift to $\widetilde{\Omega}$ and we let $\chi=(-1,0,-1)_{\xi}$ and $\chi^{\prime}=(0,0,-1)_{\xi}$. Note that $S_{\xi}=S$ forces $\xi=\det^{a}$ for some $a\in{\mathbb{Z}}$ (c.f. 5.26 later). We now have $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}=T(\mathbb{F}_{q})\times\langle\hat{\omega}^{3}\rangle$. Once again we let $V$ and $V^{\prime}$ be irreducible, finite dimensional representations of $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$ and $\chi^{\prime}$ respectively. Note that they are one dimensional because $\hat{\omega}^{3}$ is central in $\widetilde{W}$, and thus are uniquely determined by the scalars $\lambda,\lambda^{\prime}\in k^{\times}$ by which $\hat{\omega}^{3}$ acts on $V$ and $V^{\prime}$ respectively. Assume that $\lambda=\lambda^{\prime}$. We obtain two nonisomorphic, $3$-dimensional simple modules ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=(\chi^{\prime}\otimes V^{\prime})\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ of infinite projective dimension. We claim that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$.
We again have that ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$ and ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}}^{\prime})$ are projective for any face $F\subseteq\overline{C}$ of positive dimension. Applying [19, Lemma 3.4], we deduce that ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}^{\dagger}_{F}}({\mathfrak{m}})$ and ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}^{\dagger}_{F}}({\mathfrak{m}}^{\prime})$ have finite projective dimension for all such faces. By considering the Ollivier-Schneider resolution we get that the maps $\varepsilon_{\mathfrak{m}}:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\otimes_{{\mathcal{H}}^{\dagger}_{x}}{\mathcal{H}}\to{\mathfrak{m}}$ and $\varepsilon_{{\mathfrak{m}}^{\prime}}:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}}^{\prime})\otimes_{{\mathcal{H}}^{\dagger}_{x}}{\mathcal{H}}\to{\mathfrak{m}}^{\prime}$ are trivial fibrations. It again suffices to show that ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}})\cong{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}}^{\prime})$ in ${\operatorname{Ho}}({\mathcal{H}}^{\dagger}_{x})$. Note that $\Omega_{x}=\langle\omega^{3}\rangle$.

### v4

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
where $F_{1}$ is as in the proof. The second isomorphism follows from the Quillen adjunction in 5.6(ii) and the third isomorphism holds by 5.6(i) because $\chi\neq\chi^{\prime}$ for any other face in $\mathscr{F}_{0}$. Thus $\dim_{k}[\chi,\chi^{\prime}]_{\mathcal{H}_{\operatorname{aff}}}=1$ in this case. We suspect that ${\operatorname{Ho}}(\Delta)$ might be faithful in general, perhaps identifying bijectively a morphism set $[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}_{\operatorname{aff}}}}$ in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$ with the set of all $(f_{F})\in\prod_{F\subseteq\overline{C}}[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}$ such that ${\operatorname{Ho}}({\operatorname{Res}}_{{\mathcal{H}}_{F}}^{{\mathcal{H}}_{F^{\prime}}})(f_{F^{\prime}})=f_{F}$ whenever $F^{\prime}\subseteq\overline{F}$. When the group $\mathbb{G}$ has semisimple rank 1, this is in fact true. Indeed, as the complex (5.3) is a short exact sequence in that case and since ${\mathcal{H}}_{C}$ is always a semisimple algebra, we have that the augmentation again realises $\bigoplus_{F\in\mathscr{F}_{0}}{\mathcal{H}_{\operatorname{aff}}}\otimes_{{\mathcal{H}}_{F}}{\mathfrak{m}}$ as a Gorenstein projective replacement of ${\mathfrak{m}}$, for any ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$. After chasing through the adjunctions one deduces, for any ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}({\mathcal{H}_{\operatorname{aff}}})$, that the isomorphism $[{\mathfrak{m}},\mathfrak{n}]_{\mathcal{H}_{\operatorname{aff}}}\cong\bigoplus_{F\in\mathscr{F}_{0}}[{\mathfrak{m}},\mathfrak{n}]_{{\mathcal{H}}_{F}}$ obtained from the augmentation agrees with the map induced by ${\operatorname{Ho}}(\Delta)$. Finally, the compatibility condition between the $f_{F}$’s with $F\in\mathscr{F}_{0}$ and $f_{C}$ is empty here because ${\operatorname{Ho}}({\mathcal{H}}_{C})=0$.
We now study morphisms between simple supersingular ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$. Recall the parametrisation of these modules by pairs $(\chi,V)$ as in the discussion preceding 5.3.
@@
Suppose that ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=(\chi^{\prime}\otimes V^{\prime})\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ are two simple supersingular ${\mathcal{H}}$-modules satisfying ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$. Assume that ${\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}})=\infty$ and that $\chi,\chi^{\prime}$ and $\Phi$ are not as in 5.8(i). Then $\chi^{\prime}=\chi^{\tilde{\omega}}$ for some $\tilde{\omega}\in\widetilde{\Omega}$.
@@
By 2.13, we have that restriction of scalars ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}_{\operatorname{aff}}}}$ is a right Quillen functor which preserves all weak equivalences. If ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ we therefore have that ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}_{\operatorname{aff}}}})({\mathfrak{m}})$ and ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}_{\operatorname{aff}}}})({\mathfrak{m}}^{\prime})$ are also isomorphic in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$. By (5.4) we have
@@
Assume that $\chi^{\prime}\neq\chi^{\tilde{\omega}}$ for all $\tilde{\omega}\in\widetilde{\Omega}$. Since we get from the above that
@@
it then follows from 5.8(ii) that $[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{\mathcal{H}_{\operatorname{aff}}}$ has image zero under ${\operatorname{Ho}}(\Delta)$. If $f:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ denotes an isomorphism in ${\operatorname{Ho}}({\mathcal{H}})$, then this implies that ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}})(f)$ is both zero and an isomorphism for every $F\subseteq\overline{C}$. We deduce that ${\operatorname{Res}}^{{\mathcal{H}}}_{{\mathcal{H}}_{F}}({\mathfrak{m}})$, and thus ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$, is projective for every $F\subseteq\overline{C}$. Applying 5.6(iii) and 5.4 it follows that ${\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}})<\infty$, contradicting our assumptions. ∎
The conclusion of 5.19 is false when we are in the setting of 5.8(i), as is visible from the following examples.
@@
Suppose that $G=\mathrm{PGL}_{3}(\mathfrak{F})$ and assume that $p\neq 3$. Then $\Omega=\langle\omega\rangle\cong{\mathbb{Z}}/3{\mathbb{Z}}$ where $\omega$ is the image of $\left(\begin{smallmatrix}0&0&\varpi\\ 1&0&0\\ 0&1&0\end{smallmatrix}\right)$ in $W$. Similarly, let $\hat{\omega}$ denote the image of the same matrix in $\widetilde{W}$. We have $S=\{s,s^{\prime},s^{\prime\prime}\}$ and let $\chi\neq\chi^{\prime}$ be two supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ as in 5.8(i), i.e. corresponding to the pairs $(J,\xi)$ and $(J^{\prime},\xi)$ with $S_{\xi}=S$, $J=\{s,s^{\prime}\}$ and $J^{\prime}=\{s^{\prime}\}$. We will use the notation $\chi=(-1,-1,0)_{\xi}$ and $\chi^{\prime}=(0,-1,0)_{\xi}$ for these characters, the entries of the tuple corresponding to the value of the character at $T_{\hat{s}}$, $T_{\hat{s^{\prime}}}$ and $T_{\hat{s^{\prime\prime}}}$, taken in that order. Note that $S=S_{\xi}$ forces the character $\xi$ to be fixed by $\omega$ (cf. 5.29 later, viewing $\xi$ here as a character of the torus of $\mathrm{GL}_{3}(\mathbb{F}_{q})$ which is trivial on scalar matrices). Also, the element $\omega$ acts on the set $S$ by permuting its elements cyclically and correspondingly we see that $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}=T(\mathbb{F}_{q})$ and ${\mathcal{H}}_{\chi}={\mathcal{H}}_{\chi^{\prime}}={\mathcal{H}_{\operatorname{aff}}}$. If we let $V$ and $V^{\prime}$ be two irreducible finite dimensional representations of $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$ and $\chi^{\prime}$ respectively, we then automatically have $V=\chi$ and $V^{\prime}=\chi^{\prime}$. Thus we obtain two nonisomorphic simple supersingular ${\mathcal{H}}$-modules by taking ${\mathfrak{m}}=\chi\otimes_{{\mathcal{H}_{\operatorname{aff}}}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=\chi^{\prime}\otimes_{{\mathcal{H}_{\operatorname{aff}}}}{\mathcal{H}}$. Furthermore, those have infinite projective dimension (cf. 5.4). We claim that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$.
To see this, first note that (5.4) gives in this example that
(5.9) ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}{\mathfrak{m}}\cong\chi_{0}\oplus\chi_{1}\oplus\chi_{2}:=(-1,-1,0)_{\xi}\oplus(0,-1,-1)_{\xi}\oplus(-1,0,-1)_{\xi},$
and similarly that
(5.10) ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}{\mathfrak{m}}^{\prime}\cong\chi^{\prime}_{0}\oplus\chi^{\prime}_{1}\oplus\chi^{\prime}_{2}:=(0,-1,0)_{\xi}\oplus(0,0,-1)_{\xi}\oplus(-1,0,0)_{\xi}.$
@@
Next, observe that for every face $F\subseteq\overline{C}$ with $F\neq C$ the stabiliser $\Omega_{F}$ of $F$ in $\Omega$ is trivial. Thus, for such faces, ${\mathcal{H}}^{\dagger}_{F}={\mathcal{H}}_{F}$ by [27, Lemma 4.20(i)]. Let $x$ denote the vertex of $\overline{C}$ with $S_{x}=\{s^{\prime},s^{\prime\prime}\}$. By our assumption on $p$ and since all faces in $\overline{C}$ of a given dimension are $\Omega$-conjugate to each other, we obtain by applying 5.17(ii) that the augmentation $\varepsilon:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\otimes_{{\mathcal{H}}_{x}}{\mathcal{H}}\to{\mathfrak{m}}$ in the Ollivier-Schneider resolution is a trivial fibration. Here we dropped the twist by orientation $\epsilon_{x}$ character from the notation as it is trivial (since the stabiliser of $x$ automatically fixes it pointwise). Completely analogously, $\varepsilon:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}}^{\prime})\otimes_{{\mathcal{H}}_{x}}{\mathcal{H}}\to{\mathfrak{m}}^{\prime}$ is also a trivial fibration.
@@
Now, by appealing to 5.12(ii) we see that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{1})$ and ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{2})$ are projective, while ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{0})={\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{0})$ and ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{2})={\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{1})$ are not projective. By plugging all this into (5.9) and (5.10), we get a sequence of isomorphisms
${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\cong{\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{0})\oplus{\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi_{2})={\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{0})\oplus{\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{x}}(\chi^{\prime}_{1})\cong{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}}^{\prime})$
@@
Suppose that $G=\mathrm{GL}_{3}(\mathfrak{F})$. Then $\Omega=\langle\omega\rangle\cong{\mathbb{Z}}$ where $\omega$ is once again the image of $\left(\begin{smallmatrix}0&0&\varpi\\ 1&0&0\\ 0&1&0\end{smallmatrix}\right)$ in $W$. We keep the same notation as in the previous example, so we let $\hat{\omega}$ be the same lift to $\widetilde{\Omega}$ and we let $\chi=(-1,-1,0)_{\xi}$ and $\chi^{\prime}=(0,-1,0)_{\xi}$. Note that $S_{\xi}=S$ forces $\xi=\det^{a}$ for some $a\in{\mathbb{Z}}$ (cf. 5.29 later). We now have $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}=T(\mathbb{F}_{q})\times\langle\hat{\omega}^{3}\rangle$. Once again we let $V$ and $V^{\prime}$ be irreducible, finite dimensional representations of $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$ and $\chi^{\prime}$ respectively. Note that they are one dimensional because $\hat{\omega}^{3}$ is central in $\widetilde{W}$, and thus are uniquely determined by the scalars $\lambda,\lambda^{\prime}\in k^{\times}$ by which $\hat{\omega}^{3}$ acts on $V$ and $V^{\prime}$ respectively. Assume that $\lambda=\lambda^{\prime}$. We obtain two nonisomorphic, $3$-dimensional simple modules ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=(\chi^{\prime}\otimes V^{\prime})\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ of infinite projective dimension. We claim that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$.